A meter stick moving at 0.900c relative to the Earth's surface approaches an observer at rest with respect to the Earth's surface. What is the meter sticks length as measured by the observer? Qualitatively, how would the answer to that change if the observer started running toward the meter stick?
1
Expert's answer
2015-05-08T04:17:56-0400
Solution:
L' = L*sqrt( 1 - V^2/C^2) = L*sqrt(1 - 0.95*0.95) = L* sqrt(1 - 0,9025) = L* sqrt(0,0975) = L*0,312
The linear length L' of the stick moving at a speed = 0.9 C for a stationary observer will be
0.312 * L, where L is the linear size of the ship stationary relative to observer
If L = 1 m then L' = 0,312 m
Qualitatively, how would the answer to that change if the observer started running toward the meter stick?
Answer:
I think if the meter stick is stationary observer moves but the situation has not changed
meter stick will appear to the observer is shorter
and its length will be the same 0.312 m